OpenStudy (anonymous):

A parent died and left an estate to four children. One inherited 1/2 of the estate, the second inherited 1/8, and the third inherited 1/4. How much did the fourth inherit?

OpenStudy (anonymous):

0.0625

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

it is on the basis of half split. The more inheritents the more spilt.

OpenStudy (anonymous):

this question sounds nazi to me

OpenStudy (anonymous):

one minus one half is on half one half minus one quarter is one quarter one quarter minus one eighth is one eighth that is what is left

OpenStudy (anonymous):

Yeap 5/8

OpenStudy (anonymous):

\[1-\frac{1}{2}=\frac{1}{2}\] \[\frac{1}{2}-\frac{1}{4}=\frac{2}{4}-\frac{1}{4}=\frac{1}{4}\] \[\frac{1}{4}-\frac{1}{8}=\frac{2}{8}-\frac{1}{8}=\frac{1}{8}\]

OpenStudy (anonymous):

Oops...hes right cuz my answer applies to nth number of children where the number of children is i.e infinity.

OpenStudy (anonymous):

Oh ok that makes sense thank you!